# Topology of shallow-water waves on the rotating sphere

### v1

@@ 4.2 Degeneracy points of the symbol and their Chern numbers
Let us now determine the topological properties of the matrix $H$ of Expression (12). As a function of $(m,K_{\theta},\theta)$ and a 3-by-3 Hermitian matrix, $H$ has generically 3 eigenvalues which are real-valued functions defined over the parameter space $(m,K_{\theta},\theta)$. These will be referred to as eigenbands in the following, noted $\omega_{n}$ with $n=-1,0,+1$ for increasing eigenvalues. It is important to understand that these eigenvalues are not the same as the eigenfrequencies $\omega$ of the operator $\mathcal{H}$ as defined in (10).
The fields in Equations (2.1) are real-valued. As such, we have $\overline{\mathcal{H}(m)}=-\mathcal{H}(-m)$, where the overline stands for the complex conjugation of all coefficients of $\mathcal{H}$. Consequently, the spectrum of $\mathcal{H}$ is completely symmetric under the inversion $(m,\omega)\rightarrow(-m,-\omega)$. In the same way, all eigenvalues of $H$ verify
$\omega_{-n}(m,K_{\theta},\theta)=-\omega_{+n}(-m,-K_{\theta},\theta)\ ,$ (14)
for any band index $n=-1,0,+1$. Therefore, it is sufficient to consider only the positive eigenfrequencies $\omega$ of $\mathcal{H}$, and the eigenbands $n=0$ (quasi-local representation of Rossby waves) and $n=+1$ (quasi-local representation of inertia-gravity waves) of $H$. For this reason, even though a given degeneracy point of $H$ has three Chern numbers (one for each eigenband $n=-1,0,+1$), we will mostly consider the $n=+1$ one and simply refer to it as Chern number (without mentioning the band index). The characteristic polynomial of the matrix $H$ reads as
@@
The polynomial (15) has a multiple root, i.e. $H$ has a degenerated eigenvalue ($\omega_{0}=\omega_{+1}$ and/or $\omega_{0}=\omega_{-1}$), if and only if
@@
as computed in [19] for the topographic shallow-water model. The degenerated positive eigenvalue ($\omega_{0}=\omega_{+1}$) is equal to $|f|$, and the corresponding degeneracy points in parameter space are the ones such that $\beta_{\rm g}(\theta)=\pm f(\theta)$, $M=\mp|f|$ and $K_{\theta}=0$. Now the modal flow of $\mathcal{H}$ for a eigenband $n$, i.e. the number of modes gained by the Rossby (for $n=0$) or the inertia-gravity (for $n=+1$) waveband as $m$ goes from $\infty$- ∞ to +$\infty$+ ∞, is constrained by the presence of at least one of these degeneracy points in the real system, and the value of this modal flow is given by the Chern number (a negative Chern number corresponds to a net loss of modes as $m$ increases) held by this degeneracy point for the band $n$ of the symbol [1, 15, 38]. Since $f$ and $\beta_{\rm g}$ have the same sign (negative in the southern hemisphere and positive in the northern one), the problem amounts to solving $c\tan(\theta)/2R=2\Omega\sin(\theta)$, which yields
@@
  • If $\epsilon<4$, there are three latitudes at which a degeneracy between the eigenvalues of the symbol occurs, one at the equator (3-fold since all three eigenvalues of $H$ are degenerated there) and two (2-fold) degeneracy points near the poles, at opposite latitude. As $\epsilon$ increases, these get closer to the equator, where they eventually merge with the equatorial degeneracy point for $\epsilon=4$. The value of $m$ at the degeneracy points for the positive eigenbands is
@@
  • If $\epsilon\geq 4$, there is a unique degeneracy point at the equator. This can be seen as a threshold in the competition between two different $\beta$-effects, the traditional one owing to the variation of $f$ with latitude, which tends to trap waves at the equator, and the geometric one, which conversely tends to spread the eigenfunctions away from it, over the whole sphere.
Figure 9: Phase diagram of $H$, inspired by [70]. The colored curves represent the trajectory of $(\beta_{\rm g},f)$ as $\theta$ goes from $-\pi/2$ to $+\pi/2$. $H$ has a degenerated eigenvalue when $f=\pm\beta_{\rm g}$ (dashed lines), thus splitting the parameter space $(\beta_{\rm g},f)$ in 4 areas, each of which is characterised by an integer $0,-1$ or $+1$. The Chern number $C_{+1}$ (colored discs) of a degeneracy point (black dots) involving the eigenband $n=+1$ is given by their difference as $\theta$ increases past the degeneracy point. Middle: if $\epsilon<4$, the equatorial degeneracy point has $C_{+1}=+2$, same as in [1], and there are two additional mid-latitude degeneracy points with $C_{+1}=-1$. Left: as $\epsilon$ increases and reaches $4$, they merge with the equatorial degeneracy point, which still exists for $\epsilon>4$ and has zero Chern number. Right: the symbol of shallow-water waves on the unbounded $\beta$-plane is obtained by taking $\beta_{g}=0$ and $f=\beta y$ in (12), which yields the unique degeneracy point of Chern number $+2$ studied in [1].
Each of these degeneracy points can be assigned a set of Chern numbers, one for each waveband involved in the degeneracy (see Appendix 6.3). Again, two situations can be distinguished:
@@
  • For $\epsilon\geq 4$, all degeneracy points merge at the equator – and their respective Chern numbers add up – into a unique three-fold degeneracy point which is topologically neutral, i.e. of zero Chern numbers: $(C_{-1},C_{0},C_{+1})=(0,0,0)$. In the spectrum of the wave operator, this fusion manifests as a symmetrisation of the eigenvalues and eigenmodes between positive and negative $m$, whereas the east-west asymmetry of the spectrum is appreciable for small $\epsilon$, when the degeneracy points are separated.
In conclusion, the metric term $\beta_{g}$ generates systematic degeneracy points whose Chern numbers compensate that of the equatorial degeneracy point. For small $\epsilon$, these degeneracy points are located near the poles and bear the Chern number $-1$ for the eigenband $n=+1$, and for $\epsilon\geq 4$ there is a unique degeneracy point at the equator whose Chern numbers are 00. The eigenbands of $H$ thus have zero total Chern numbers, which confirms the nullity of the net modal flow in the spectrum of shallow-water waves on the rotating sphere, in virtue of the index theorem.
@@ 4.3 Collapse of the spectral flow
  • Although the index theorem applies individually for each degeneracy point of the symbol, the resulting spectral flows are visible only in a semi-classical limit [15], i.e. as long as the transiting modes are well-separated in phase space [71]. In the case of shallow-water waves on the rotating sphere, the semi-classical limit in question corresponds to $\epsilon\ll 1$. Indeed, the equatorial waves spread up to the poles even for moderate values of $\epsilon$ ($\epsilon\sim 1$), which means that the modes corresponding to the opposite spectral flows overlap and hybridise, thus manifesting an avoided crossing between their respective spectral branches (see e.g. the cases studied in [72] and [38], part 3.5.1), which is precisely why the spectral flow of eigenfrequencies collapses as $\epsilon$ increases.
@@ 5 Concluding remarks
The goal of this work is to investigate the topological properties of the shallow-water spectrum on the rotating sphere. In particular, we aimed at answering the following questions: how do the results of [1], which were established using the $\beta$-plane approximation, extend to the spherical case? What does topology tell us about the transition between the Matsuno and Margules limits? Both regimes have been extensively investigated, and the corresponding solutions can be well-approximated, however there is no analytic solution for arbitrary values of $\epsilon$. In the Matsuno limit ($\epsilon\ll 1$ or large Lamb parameter), there seems to be a continuous spectral connection between the Rossby and inertia-gravity modes, which is embodied by the Yanai and Kelvin modes, whose frequencies cross the gap as $m$ increases (see Figure 1). This spectral feature is reminiscent of the Matsuno spectrum on the unbounded $\beta$-plane, for which the connection with topological invariant is well-established [1, 15, 16, 69]. In the Margules limit (large $\epsilon$ or small Lamb parameter), this gapless connection progressively breaks down. All the inertia-gravity modes plus the positive-phase-speed Yanai (with $m\geq 0$) and Kelvin modes (with $m\geq 1$) move up to values $\omega=\mathcal{O}(c/R)$, while the largest Yanai eigenfrequency ($m=-1$) remains smaller than $\Omega$ (see Equation (7) with $n=0$ and Figure 4), thus separating the spectrum into two distinct wavebands with an open frequency gap.
@@
For small $\epsilon$, the eigenfrequencies of shallow-water waves on the sphere and those of the Matsuno spectrum on the unbounded $\beta$-plane are nearly identical, however the $\tilde{v}$ component of the eigenmodes of positive (eastward) phase speed has two more zeros located at opposite latitudes. The latter being close to the poles, they are not captured by the $\beta$-plane approximation. Eigenmodes on the sphere have been extensively studied, in particular the Yanai modes [73, 43]. Nevertheless, to our knowledge, the existence of these additional zeros is usually not discussed in the geophysical literature. Yet they are consistent with the evolution of the shallow-water spectrum as $\epsilon$ varies continuously. Indeed, consider an eigenmode with fixed wave number $m$ and frequency $\omega$ for a large value of $\epsilon$ (Margules limit). The number $p$ of zeros of $\tilde{v}(\theta)$ is established by the classification given in 2.2. The inertia-gravity eigenmodes are essentially those of the Laplacian on the sphere, which are insensitive to the sign of $m$. As $\epsilon$ continuously decreases, the effect of rotation becomes stronger and the spectrum progressively loses this east-west symmetry. For the mode picked in the large $\epsilon$ regime, its frequency $\omega$ and the eigenfunction $\tilde{v}$ change as well in a continuous manner. However, the number $p$ cannot change as the structure of $\tilde{v}$ is governed by a generalised Sturm-Liouville problem (see the arguments of [59]), whose eigenvalue equation is the spherical generalisation of Matsuno’s equation (Equation (3.3) in [34] is given for the perturbation of potential vorticity, which is proportional to $\cos{(\theta)}\tilde{v}$). When decreasing $\epsilon$, the frequencies of the eastward modes decrease faster than the westward ones, leading to a dislocation (or $+2$ branch jump) in the spectrum. This dislocation eventually results in the spectral flow of Yanai and Kelvin modes in the small $\epsilon$ (Matsuno) limit. In the Matsuno limit on the sphere, this spectral flow is thus obtained as a continuous deformation of the Margules spectrum with no spectral flow. It arises from the isolated $+2$ topological charge at the equator, whereas the opposite topological charges near the poles generate invisible modes with opposite spectral flow, which manifest in the spectrum through the change in number of $\tilde{v}$ zeros of Yanai modes as $m$ changes sign. In other words, this spectral flow is not a modal flow, since the westward and eastward Yanai modes do not constitute a single modal branch. In contrast, the shallow-water spectrum on the unbounded $\beta$-plane arises from a single non-zero topological charge, because there is no such way to continuously deform it into a spectrum with no modal flow, as the parameters $\beta$ and $c$ can be absorbed in the definition of length and time units, and thus do not affect the shape of the spectrum discussed in 3.1. Conversely, for waves on the sphere, both $R$ and the Rossby radius of deformation $c/2\Omega$ deeply affect the shape of the solutions, and their ratio define the effective equatorial trapping of waves.
@@
We then highlighted furthermore this distinction between the unbounded $\beta$-plane and the sphere by investigating the topology of the symbol of the shallow-water model on the sphere. We showed that the shallow-water wavebands on the sphere have zero total Chern numbers, in contrast with the wavebands on the unbounded $\beta$-plane. Consequently, in virtue of the index theorem, there is no net modal flow of modes in the spectrum. Precisely, we found that when $\epsilon$ becomes higher than $4$, several degeneracy points merge into a unique, topologically neutral degeneracy point at the equator. This manifests in the shallow-water spectrum as the loss of east-west asymmetry and the complete collapse of the spectral flow of equatorial Yanai and Kelvin waves. However, for $\epsilon<4$, the symbol of the wave operator on the sphere has several charged degeneracy points: an equatorial one with non-zero Chern numbers, which is reminiscent of the unbounded $\beta$-plane [1], plus two others (considering the positive eigenbands $n=0,+1$) with opposite charges. As discussed in 4.3, the spectral footprint of the former are the Yanai and Kelvin waves, whose transiting eigenfrequencies are characteristic of the east-west asymmetry at small $\epsilon$, and that of the latter is the $+2$ modal jump, i.e. the fact that opposite topological charges of the symbol lead, for small $\epsilon$, to a spectral flow that is not a modal flow.
@@
This work joins recent efforts toward the comprehension of how topology applies in the spectral properties of continuous physical systems on curved surfaces [26, 74, 75, 64], a subject that still lacks a unified framework. In that sense, we wish to stress the importance of the Weyl symbol and the spectral parameter. With this study, we have shown that the shallow-water model on the rotating sphere has a counter-intuitive topology that is either neutral (all Chern numbers are null for $\epsilon\geq 4$) or polar in the sense that it has several degeneracy points of non-zero Chern numbers which sum to zero (for $\epsilon<4$), whereas its $\beta$-plane counterpart in unbounded flat geometry has a unique, topologically charged degeneracy point. Besides, we exhibited a peculiar situation, in which different degeneracy points (with different multiplicity) merge and yield a unique degeneracy point which is topologically neutral, yet without the gap opening afterward. This situation is not usual in topological physics. Further investigation is necessary to better understand the spectral properties of continuous systems on curved surfaces, and the manifestation of topology in their dynamics.

### v2

@@ 4.2 Degeneracy points of the symbol and their Chern numbers
Let us now determine the topological properties of the matrix $H$ of Expression (13). As a function of $(m,K_{\theta},\theta)$ and a 3-by-3 Hermitian matrix, $H$ has generically 3 eigenvalues which are real-valued functions defined over the parameter space $(m,K_{\theta},\theta)$. These will be referred to as eigenbands in the following, noted $\lambda_{n}$ with $n=-1,0,+1$ for increasing eigenvalues. It is important to understand that these eigenvalues are not the same as the frequencies $\omega$, i.e. the eigenvalues of the operator $\mathcal{H}$ as defined in (11).
The fields in Equations (3) are real-valued. As such, we have $\overline{\mathcal{H}(m)}=-\mathcal{H}(-m)$, where the overline stands for the complex conjugation of all coefficients of $\mathcal{H}$. Consequently, the spectrum of $\mathcal{H}$ is completely symmetric under the inversion $(m,\omega)\rightarrow(-m,-\omega)$. In the same way, all eigenvalues of $H$ verify
$\lambda_{-n}(m,K_{\theta},\theta)=-\lambda_{+n}(-m,-K_{\theta},\theta)\ ,$ (14)
for any band index $n=-1,0,+1$. Therefore, it is sufficient to consider only the positive frequencies $\omega$ of $\mathcal{H}$, and the eigenbands $n=0$ (quasi-local representation of Rossby waves) and $n=+1$ (quasi-local representation of inertia-gravity waves) of $H$. For this reason, even though a given degeneracy point of $H$ has three Chern numbers (one for each eigenband $n=-1,0,+1$), we will mostly consider the $n=+1$ one and simply refer to it as Chern number (without mentioning the band index). The characteristic polynomial of the matrix $H$ reads as
@@
The polynomial (15) has a multiple root, i.e. $H$ has a degenerated eigenvalue ($\lambda_{0}=\lambda_{+1}$ and/or $\lambda_{0}=\lambda_{-1}$), if and only if
@@
as computed by [27] for the topographic shallow-water model. The degenerated positive eigenvalue ($\lambda_{0}=\lambda_{+1}$) is equal to $|f|$, and the corresponding degeneracy points in parameter space are the ones such that $\beta_{\rm g}(\theta)=\pm f(\theta)$, $M=\mp|f|$ and $K_{\theta}=0$. Now the modal flow of $\mathcal{H}$ for an eigenband $n$, i.e. the number of modes gained by the Rossby (for $n=0$) or the inertia-gravity (for $n=+1$) waveband as $m$ goes from $\infty$- ∞ to +$\infty$+ ∞, is constrained by the presence of at least one of these degeneracy points in the real system, and the value of this modal flow is given by the Chern number (a negative Chern number corresponds to a net loss of modes as $m$ increases) held by this degeneracy point for the band $n$ of the symbol [1, 23, 37]. Since $f$ and $\beta_{\rm g}$ have the same sign (negative in the southern hemisphere and positive in the northern one), the problem amounts to solving $c\tan(\theta)/2R=2\Omega\sin(\theta)$, which yields
@@
  • If $\epsilon<4$, there are three latitudes at which a degeneracy between the eigenvalues of the symbol occurs, one at the equator (3-fold since all three eigenvalues of $H$ are degenerated there) and two (2-fold) degeneracy points near the poles, at opposite latitude. As $\epsilon$ increases, these get closer to the equator, where they eventually merge with the equatorial degeneracy point for $\epsilon=4$. The value of $m$ at the non-equatorial degeneracy points for the positive eigenbands is
@@
  • If $\epsilon\geq 4$, there is a unique degeneracy point at the equator. This can be seen as a threshold in the competition between two different $\beta$-effects, the traditional one owing to the variation of $f$ with latitude, which tends to trap waves at the equator, and the geometric one, which conversely tends to spread the wave functions away from it, over the whole sphere.
Figure 9: Phase diagram of $H$, inspired by [69]. The colored curves represent the trajectory of $(\beta_{\rm g},f)$ as $\theta$ goes from $-\pi/2$ to $+\pi/2$. $H$ has a degenerated eigenvalue when $f=\pm\beta_{\rm g}$ (dashed lines), thus splitting the parameter space $(\beta_{\rm g},f)$ in 4 areas, each of which is characterised by an integer $0,-1$ or $+1$. The Chern number $C_{+1}$ (colored discs) of a degeneracy point (black dots) involving the eigenband $n=+1$ is given by their difference as $\theta$ increases past the degeneracy point. Middle: if $\epsilon<4$, the equatorial degeneracy point has $C_{+1}=+2$, same as in [1], and there are two additional mid-latitude degeneracy points with $C_{+1}=-1$. Left: as $\epsilon$ increases and reaches $4$, they merge with the equatorial degeneracy point, which still exists for $\epsilon>4$ and has zero Chern number. Right: the symbol of shallow-water waves on the unbounded $\beta$-plane is obtained by taking $\beta_{\rm g}=0$ and $f=\beta y$ in (13), which yields the unique degeneracy point of Chern number $+2$ studied by [1].
Each of these degeneracy points can be assigned a set of Chern numbers, one for each eigenband involved in the degeneracy (see Appendix 6.4). Again, two situations can be distinguished:
@@
  • For $\epsilon\geq 4$, all degeneracy points merge at the equator – and their respective Chern numbers add up – into a unique three-fold degeneracy point which is topologically neutral, i.e. of zero Chern numbers: $(C_{-1},C_{0},C_{+1})=(0,0,0)$. In the spectrum of the wave operator, this fusion manifests as a symmetrisation of the modes between positive and negative $m$, whereas the east-west asymmetry of the spectrum is appreciable for small $\epsilon$, when the degeneracy points are separated.
In conclusion, the metric term $\beta_{\rm g}$ generates systematic degeneracy points whose Chern numbers compensate that of the equatorial degeneracy point. For small $\epsilon$, these degeneracy points are located near the poles and bear the Chern number $-1$ for the eigenband $n=+1$, and for $\epsilon\geq 4$ there is a unique degeneracy point at the equator whose Chern numbers are 00. The eigenbands of $H$ thus have zero total Chern numbers, which confirms that the net number of modes gained by the inertia-gravity waveband as $m$ is swept from $\infty$- ∞ to +$\infty$+ ∞ is zero, in virtue of the index theorem. As explained in section 3, there is a direct correspondence between this mode imbalance and the evolution of the number of zeros of $\tilde{v}$ for waves on the unbounded $\beta$-plane. While the same correspondence can only be inferred by extrapolation for waves on the rotating sphere, we showed here that the absence of modal flow is in agreement with the predictions of topology. For the unbounded $\beta$-plane, [62] provides a formal proof of the equality between the modal flow (i.e. the difference in number of positive-frequency inertia-gravity modes between the two limits $k\rightarrow+\infty$ and $k\rightarrow-\infty$) and the Chern number of the symbol’s degeneracy point at the equator, based on geometric WKB analysis of the waves in both limits. This study, which provides an alternative point of view to the index theorem, could also be extended here to demonstrate the nullity of this mode imbalance on the sphere, with the wave number $m$ instead of $k$.
@@ 4.3 Collapse of the spectral flow
  • Although the index theorem applies individually for each degeneracy point of the symbol, the resulting spectral flows are visible only in a semi-classical limit [23], i.e. as long as the transiting modes are well-separated in phase space [70]. In the case of shallow-water waves on the rotating sphere, the semi-classical limit in question corresponds to $\epsilon\ll 1$. Indeed, the equatorial waves spread up to the poles even for moderate values of $\epsilon$ ($\epsilon\sim 1$), which means that the modes corresponding to the opposite spectral flows overlap and hybridise, thus manifesting an avoided crossing between their respective spectral branches (see e.g. the cases studied by [71] and [37], part 3.5.1), which is precisely why the spectral flow of frequencies collapses as $\epsilon$ increases.
@@ 5 Concluding remarks
The goal of this work is to investigate the topological properties of the shallow-water spectrum on the rotating sphere. In particular, we aimed at answering the following questions: how do the results of [1], which were established using the $\beta$-plane approximation, extend to the spherical case? What does topology tell us about the transition between the Matsuno and Margules limits? Both regimes have been extensively investigated, and the corresponding solutions can be well-approximated, however there is no analytic solution for arbitrary values of $\epsilon$. In the Matsuno limit ($\epsilon\ll 1$ or large Lamb parameter), there seems to be a continuous spectral connection between the Rossby and inertia-gravity modes, which is embodied by the Yanai and Kelvin modes, whose frequencies cross the gap as $m$ increases (see Figure 1). This spectral feature is reminiscent of the Matsuno spectrum on the unbounded $\beta$-plane, for which the connection with topological invariant is well-established [1, 23, 24, 62]. In the Margules limit (large $\epsilon$ or small Lamb parameter), this gapless connection progressively breaks down. All the inertia-gravity modes plus the positive-phase-speed Yanai (with $m\geq 0$) and Kelvin modes (with $m\geq 1$) move up to values $\omega=\mathcal{O}(c/R)$, while the largest Yanai frequency ($m=-1$) remains smaller than $\Omega$ (see Equation (8) with $n=0$ and Figure 4), thus separating the spectrum into two distinct wavebands with an open frequency gap.
@@
For small $\epsilon$, the frequencies of shallow-water waves on the sphere and those of the Matsuno spectrum on the unbounded $\beta$-plane are nearly identical, however the $\tilde{v}$ component of the modes of positive (eastward) phase speed has two more zeros located at opposite latitudes. The latter being close to the poles, they are not captured by the $\beta$-plane approximation. Modes on the sphere have been extensively studied, in particular the Yanai modes [72, 42]. Nevertheless, to our knowledge, the existence of these additional zeros is usually not discussed in the geophysical literature. Yet they are consistent with the evolution of the shallow-water spectrum as $\epsilon$ varies continuously. To explain why, let us consider a mode with wave number $m$ and frequency $\omega$, for a large value of $\epsilon$ (Margules limit). The number $p$ of zeros of $\tilde{v}(\theta)$ is established by the classification given in 2.2. The inertia-gravity modes are essentially those of the Laplacian on the sphere, which are insensitive to the sign of $m$. As $\epsilon$ continuously decreases, the effect of rotation becomes stronger and the spectrum progressively loses this east-west symmetry. For the mode picked in the large $\epsilon$ regime, its frequency $\omega$ and the wave function $\tilde{v}$ change as well in a continuous manner. However, the number $p$ cannot change as the structure of $\tilde{v}$ is governed by a generalised Sturm-Liouville problem (see the arguments of [58]), whose eigenvalue equation is the spherical generalisation of Matsuno’s equation (Equation (3.3) of [19] is given for the perturbation of potential vorticity, which is proportional to $\cos{(\theta)}\tilde{v}$). When decreasing $\epsilon$, the frequencies of the eastward modes decrease faster than the westward ones, leading to a dislocation (or $+2$ branch jump) in the spectrum. This dislocation eventually results in the spectral flow of Yanai and Kelvin modes in the small $\epsilon$ (Matsuno) limit. In the Matsuno limit on the sphere, this spectral flow is thus obtained as a continuous deformation of the Margules spectrum with no spectral flow. It arises from the isolated $+2$ topological charge at the equator, whereas the opposite topological charges near the poles generate invisible modes with opposite spectral flow, which manifest in the spectrum through the change in number of $\tilde{v}$ zeros of Yanai modes as $m$ changes sign. In other words, this spectral flow is not a modal flow, since the westward and eastward Yanai modes do not constitute a single modal branch. In contrast, the shallow-water spectrum on the unbounded $\beta$-plane arises from a single non-zero topological charge, because there is no such way to continuously deform it into a spectrum with no modal flow, as the parameters $\beta$ and $c$ can be absorbed in the definition of length and time units, and thus do not affect the shape of the spectrum discussed in 3.1. Conversely, for waves on the sphere, both $R$ and the Rossby radius of deformation $c/2\Omega$ deeply affect the shape of the solutions, and their ratio define the effective equatorial trapping of waves.
@@
We then highlighted furthermore this distinction between the unbounded $\beta$-plane and the sphere by investigating the topology of the symbol of the shallow-water model on the sphere. We showed that the shallow-water eigenbands on the sphere have zero total Chern numbers, in contrast with the eigenbands on the unbounded $\beta$-plane. Consequently, in virtue of the index theorem, there is no net modal flow of modes in the spectrum. Precisely, we found that when $\epsilon$ becomes higher than $4$, several degeneracy points merge into a unique, topologically neutral degeneracy point at the equator. This manifests in the shallow-water spectrum as the loss of east-west asymmetry and the complete collapse of the spectral flow of equatorial Yanai and Kelvin waves. However, for $\epsilon<4$, the symbol of the wave operator on the sphere has several charged degeneracy points: an equatorial one with non-zero Chern numbers, which is reminiscent of the unbounded $\beta$-plane [1], plus two others (considering the positive eigenbands $n=0,+1$) with opposite charges. As discussed in 4.3, the spectral footprint of the former are the Yanai and Kelvin waves, whose transiting frequencies are characteristic of the east-west asymmetry at small $\epsilon$, and that of the latter is the $+2$ modal jump, i.e. the fact that opposite topological charges of the symbol lead, for small $\epsilon$, to a spectral flow that is not a modal flow.
@@
This work joins recent efforts toward the comprehension of how topology applies in the spectral properties of continuous physical systems on curved surfaces [34, 73, 74, 75, 67], a subject that still lacks a unified framework. In that sense, we wish to stress the importance of the Weyl symbol and the spectral parameter. With this study, we have shown that the shallow-water model on the rotating sphere has a counter-intuitive topology that is either neutral (all Chern numbers are null for $\epsilon\geq 4$) or polar in the sense that it has several degeneracy points of non-zero Chern numbers which sum to zero (for $\epsilon<4$), whereas its $\beta$-plane counterpart in unbounded flat geometry has a unique, topologically charged degeneracy point. Besides, we exhibited a peculiar situation, in which different degeneracy points (with different multiplicity) merge and yield a unique degeneracy point which is topologically neutral, yet without the gap opening afterward. This situation is not usual in topological physics. Further investigation is necessary to better understand the spectral properties of continuous systems on curved surfaces, and the manifestation of topology in their dynamics.